Analyze a CSR Capital Investment Proposal for Den-Tex Company Den-Tex Company is evaluating a proposal to replace its HID (hiHello, I am having issues with this question. Please do show work, Thank you very much!
Answer #1
a. Investment cost=Life time investment cost*Number of fixtures=500*700=$ 350000
b. Annual utility cost savings:
Kilowatt-hour saved per fixture per day:
Kilowatt
HID Kilowatt hour consumption per fixture (500/1000) 0.5
Less: LED Kilowatt hour consumption per fixture (300/1000) 0.3
Kilowatt-hour saved per fixture per day: 0.2
Kilowatt-hour saved per year=Kilowatt-hour saved per fixture per day*Operating hours per day*Operating days per year*Number of fixtures=0.2*10*300*700=420000 Kilowatt-hours
Annual utility cost savings=Kilowatt-hours saved per year*Metered utility rate per kilowatt-hour=420000*0.11=$ 46200
c. Net present value=Present value of annual utility cost savings-Investment cost
Present value of annual utility cost savings=Annual utility cost savings*Discount factor at 8% for 15 years=46200*8.55948=$ 395448
Net present value=395448-350000=$ 45448
The proposal should be accepted since the net present value is positive
I appreciate your ratings